OpenStudy (anonymous):

um what...? Cadmium and zinc are used in a cell to produce cadmium metal. Using the standard reduction potentials shown, calculate the cell potential and identify the type of cell. Cd2+(aq) + 2e– Cd(s) Eº = –0.40 V Zn2+(aq) + 2e– Zn(s) Eº = –0.76 V

OpenStudy (anonymous):

@jdoe0001 @thomaster @Australopithecus

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

This sentence "Cadmium and zinc are used in a cell to produce cadmium metal." implies that \(Cd^{2+}\) will oxidize \(Zn_{(s)}\) to produce \(Cd_{(s)}\). You have to add up the individual potentials. The net equation will look like this: \(\sf Cd^{2+}+Zn_{(s)}→Zn^{2+}+Cd_{(s)}\) So the individual equations must be: \(\sf Cd^{2+} + 2e^– → Cd_{(s)} ~~~ \xi^o = –0.40 V\) Cathode \( \sf Zn(s) →2e^– + Zn^{2+} ~~~ \xi^o = \color{red}+0.76 V\) Anode \(ξ^o_{cell}=ξ^o_{cathode}+ξ^o_{anode}\) \(ξ^o_{cell}=–0.40 V+0.76 V=0.36V\)

OpenStudy (anonymous):

The answer choices are also foreign to me though A. +0.36 V; galvanic B. –0.36 V; electrolytic C. –1.16 V; electrolytic D. +1.16 V; galvanic

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

it's A. The potential is +0.36 V, and this is a galvanic cell because it flows spontaneously. If the cell potential was less than 0 V, then it would be electrolytic, which are non-spontaneous. This means that it would need an external source of power to drive the cell. The spontaneity of a cell can be assessed with: \(\Delta G^o=-nF\xi^o \), where F is faraday's constant and n represents the moles of electrons.